please tell me the answer and process. Thank you!!

oure driving down the highway late one night at 30 m/s when you spot a deer in brakes is 0.35s and the maximum deceleration of your car is 8.0 m/s. the road some distance in front of you. Your reaction time before stepping on the a) What is the total distance covered by your car from the time you first see the (8pts) deer to when you are at a complete stop? b) What is the total time it takes for you to stop? (4pts) 4. In a 5.00 km race, one runner runs at a steady 11.0 km/h and another runs at 16.0 km/h. a) How far from the finish line is the slower runner when the faster runner finishes the race? (4pts) b) How long does the faster runner have to wait at the finish line to see the slower runner cross? (2pts)
Answer #1

Distance covered before applying brakes, d1 = 30 x 0.35 = 10.5 m

To calculate distance covered after applying brakes, we will use kinematic equation,

V^2 = u^2 + 2as

0 = 30^2 - 2(8 x d2)

Distance, d2 = 56.25 m

So, total distance is 56.25 + 10.5 = 66.75 m

B) total time= 0.35+ 30/8= 4.1 s

========

Plz post other problems seperately..

Comment in case any doubt.. Goodluck